# OpenRubrics: Towards Scalable Synthetic Rubric Generation for Reward Modeling and LLM Alignment

### v2

@@ Abstract
Reward modeling lies at the core of reinforcement learning from human feedback (RLHF), yet most existing reward models rely on scalar or pairwise judgments that fail to capture the multifaceted nature of human preferences. Recent studies have explored rubrics-as-rewards (RaR) that uses structured natural language criteria that capture multiple dimensions of response quality. However, producing rubrics that are both reliable and scalable remains a key challenge. In this work, we introduce OpenRubrics, a diverse, large-scale collection of (prompt, rubric) pairs for training rubric-generation and rubric-based reward models. To elicit discriminative and comprehensive evaluation signals, we introduce Contrastive Rubric Generation (CRG), which derives both hard rules (explicit constraints) and principles (implicit qualities) by contrasting preferred and rejected responses. We further improve reliability by enforcing preference–label consistency via rejection sampling to remove noisy rubrics. Across multiple reward-modeling benchmarks, our rubric-based reward model, Rubric-RM, surpasses strong size-matched baselines by 6.8%. These gains transfer to policy models on instruction-following and biomedical benchmarks. Our results show that rubrics provide scalable alignment signals that narrow the gap between costly human evaluation and automated reward modeling, enabling a new principle-driven paradigm for LLM alignment.
@@ 4.2 Rubrics Synthesis
Contrastive Rubric Generation. Given a preference dataset $\mathcal{D}=\left\{\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-}\right)\right\}_{i=1}^{N},$ where $x_{i}$ is the prompt and $\hat{y}_{i}^{+}$ and $\hat{y}_{i}^{-}$ denote the preferred and displeased responses, respectively. Our objective is to infer rubrics $\mathcal{R}\left(x_{i}\right)$ that capture qualities a good response should satisfy and the criteria by which one response is preferred over another, using $\hat{y}_{i}^{+}$ and $\hat{y}_{i}^{-}$ as guidance. Formally, we prompt an instruction-tuned LLM $h_{\psi}$ as $\mathcal{R}\left(x_{i}\right)\sim h_{\psi}\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\ell_{i}\right),$ where $\ell_{i}$ is the preference label. The generator is asked to produce a set of discriminative evaluation criteria $\mathcal{R}\left(x_{i}\right)=\left\{c_{i,1},\ldots,c_{i,k_{i}}\right\}$, where each $c_{i,j}$ describes a specific aspect. This setting encourages the model to discover rubric dimensions that are both task-sensitive and preference-aligned.
Rubric Filtering with Preference-label Consistency. Not all generated rubrics faithfully capture the human preference signal. To ensure reliability, we conduct a consistency-based filtering step via prompting the LLM $h_{\psi}$ again. For each triplet ($x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-}$), we concatenate the full rubric $\mathcal{R}\left(x_{i}\right)$ into the context and ask the model to predict which response better aligns with the rubric $\hat{l}_{i}=h_{\psi}\left(x_{i},\mathcal{R}\left(x_{i}\right),\hat{y}_{i}^{+},\hat{y}_{i}^{-}\right),$ where $\hat{l}_{i}=(\hat{r}_{i},\hat{\ell}_{i})$ denotes the prediction rationale and the predicted preference, respectively. For pairwise judging, we retain only those rubrics that lead to predictions consistent with the original human label $\ell_{i}$:
@@
For listwise judging settings, where a single prompt is associated with multiple candidate responses, we decompose the task into multiple pairwise comparisons. In this case, we retain a rubric if it yields preference predictions consistent with human labels in at least $50\%$ of the induced pairwise comparisons, allowing robustness to occasional ambiguity while preserving discriminative power.
This yields a collection of high-quality rubrics that are both interpretable and empirically consistent with human preferences. The final rubrics-conditioned preference dataset combines prompts, paired responses, and their verified rubrics as $\mathcal{D}_{\text{rubric }}=\left\{\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i})\right)\right\}_{i=1}^{M}.$
@@ B.4 Prompts
Prompt for General Domain Judge Generation (Rubric-RM)
Prompt for Medical Domain Judge Generation (Rubric-RM)

### v3

@@ Abstract
Reward modeling lies at the core of reinforcement learning from human feedback (RLHF), yet most existing reward models rely on scalar or pairwise judgments that fail to capture the multifaceted nature of human preferences. Recent studies have explored rubrics-as-rewards (RaR) that uses structured criteria to capture multiple dimensions of response quality. However, producing rubrics that are both reliable and scalable remains a key challenge. In this work, we introduce OpenRubrics, a diverse, large-scale collection of (prompt, rubric) pairs for training rubric-generation and rubric-based reward models. To elicit discriminative and comprehensive evaluation signals, we introduce Contrastive Rubric Generation (CRG), which derives both hard rules (explicit constraints) and principles (implicit qualities) by contrasting preferred and rejected responses. We further remove noisy rubrics via preserving preference–label consistency. Across multiple reward-modeling benchmarks, our rubric-based reward model, Rubric-RM, surpasses strong size-matched baselines by 8.4%. These gains transfer to policy models on instruction-following and biomedical benchmarks.
@@ 4.2 Rubrics Synthesis
Contrastive Rubric Generation. Given a preference dataset $\mathcal{D}=\left\{\left(x_{i},\{\hat{y}_{i,m}\}_{m=1}^{M_{i}},\{\ell_{i,m}\}_{m=1}^{M_{i}}\right)\right\}_{i=1}^{N},$ where $x_{i}$ is the prompt, $\{\hat{y}_{i,m}\}_{m=1}^{M_{i}}$ and $\{\ell_{i,m}\}_{m=1}^{M_{i}}$ denotes a list of candidate responses and their corresponding preference signals, respectively. The preference signal $\ell_{i,m}$ can be a real-valued score when available, or a binary chosen–rejected label. We unify both cases by constructing a strictly ordered list in descending preference: $\hat{y}_{i,1}\succ\hat{y}_{i,2}\succ\@cdots\succ\hat{y}_{i,M_{i}},$ where the order is induced by $\ell$. In particular, when only chosen–rejected labels are available, we simply set $M_{i}=2$ with $\hat{y}_{i,1}$ as chosen and $\hat{y}_{i,2}$ as rejected. Our objective is to infer rubrics $\mathcal{R}(x_{i})$ that capture qualities a good response should satisfy and the criteria that explain preference differences across the list, leveraging rich and fine-grain supervision from the comparisons. Formally, we prompt an instruction-tuned LLM $h_{\psi}$ as $\mathcal{R}(x_{i})\sim h_{\psi}\!\left(x_{i},\{\hat{y}_{i,m}\}_{m=1}^{M_{i}},\{\ell_{i,m}\}_{m=1}^{M_{i}}\right).$ The generator is asked to produce a set of discriminative evaluation criteria $\mathcal{R}(x_{i})=\left\{c_{i,1},\ldots,c_{i,k_{i}}\right\}$, where each $c_{i,j}$ describes a specific aspect and is expected to distinguish higher-preference responses from lower-preference ones. This listwise setting encourages the model to discover rubric dimensions that are both task-sensitive and preference-aligned, while exploiting any available fine-grained preference information.
Rubric Filtering with Preference-label Consistency. Not all generated rubrics faithfully capture the human preference signal. To ensure reliability, we perform a consistency-based filtering step by prompting the same instruction-tuned LLM $h_{\psi}$ again. Specifically, for each prompt $x_{i}$ with an ordered response list $\{\hat{y}_{i,m}\}_{m=1}^{M_{i}}$, we construct a set of induced pairwise comparisons $\mathcal{P}_{i}=\{(a,b)\mid 1\le a<b\le M_{i}\},$ where the human preference label for each pair is $\ell_{i,(a,b)}=1$ (i.e., $\hat{y}_{i,a}\succ\hat{y}_{i,b}$). When the original data only provides chosen–rejected labels, we have $M_{i}=2$ and thus $|\mathcal{P}_{i}|=1$. Next, we concatenate the full rubric $\mathcal{R}(x_{i})$ into the context and ask the model to judge each induced pair:
$\hat{l}_{i,(a,b)}=h_{\psi}\!\left(x_{i},\mathcal{R}(x_{i}),\hat{y}_{i,a},\hat{y}_{i,b}\right),$ (1)
where $\hat{l}_{i,(a,b)}=(\hat{r}_{i,(a,b)},\hat{\ell}_{i,(a,b)})$ denotes the prediction rationale and the predicted preference, respectively. A rubric is considered reliable for prompt $x_{i}$ only if it yields preference predictions consistent with human labels on the induced pair set:
@@
where $\tau$ is a threshold (we use $\tau=0.5$). Finally, for each induced pair $(a,b)\in\mathcal{P}_{i}$, we retain the rubric-conditioned training instance only if (i) the rubric passes the group-level verification ($\mathrm{Acc}_{i}\ge\tau$), and (ii) the current pairwise prediction is correct ($\hat{\ell}_{i,(a,b)}=\ell_{i,(a,b)}$). Formally,
$\mathcal{R}^{*}_{i,(a,b)}=\begin{cases}\mathcal{R}(x_{i}),&\text{if}\quad\mathrm{Acc}_{i}\ge\tau\quad\text{and}\quad\hat{\ell}_{i,(a,b)}=\ell_{i,(a,b)},\\ \emptyset,&\text{otherwise.}\end{cases}$ (3)
This yields a collection of high-quality rubrics that are both interpretable and empirically consistent with human preferences. The final rubrics-conditioned preference dataset is constructed as a pairwise one
$\mathcal{D}_{\text{rubric}}$ $=\left\{\left(x_{i},\hat{y}_{i,a},\hat{y}_{i,b},\ell_{i,(a,b)},\mathcal{R}^{*}_{i,(a,b)}\right)\ \middle|\ i\in[N],\ (a,b)\in\mathcal{P}_{i},\ \mathcal{R}^{*}_{i,(a,b)}\neq\emptyset\right\}$ (4)
$\triangleq\{(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i}))\}_{i=1}^{M}.$ (5)
Namely, we flatten and re-index all retained pairs to obtain the pairwise dataset, where new instance corresponds to one retained $(i,(a,b))$ with $\hat{y}_{i}^{+}=\hat{y}_{i,a}$ and $\hat{y}_{i}^{-}=\hat{y}_{i,b}$.